Find the area of the shaded region. The graph depicts the standard normal distribution with mean 0 and standard deviation 1. Z=0.65 Statistics
Answer #1

From standard normal tables, we have:

P(Z < 0.65) = 0.7422

This is represented on graph as:

Also from standard normal tables, we have:

P( Z > 0.65) = 1 - P(Z < 0.65) = 1 - 0.7422 = 0.2578

This is represented on graph as: